# Hard-Constrained Neural Networks with Universal Approximation Guarantees

### v1

@@ Abstract
Incorporating prior knowledge or specifications of input-output relationships into machine learning models has gained significant attention, as it enhances generalization from limited data and leads to conforming outputs. However, most existing approaches use soft constraints by penalizing violations through regularization, which offers no guarantee of constraint satisfaction—an essential requirement in safety-critical applications. On the other hand, imposing hard constraints on neural networks may hinder their representational power, adversely affecting performance. To address this, we propose HardNet, a practical framework for constructing neural networks that inherently satisfy hard constraints without sacrificing model capacity. Specifically, we encode affine and convex hard constraints, dependent on both inputs and outputs, by appending a differentiable projection layer to the network’s output. This architecture allows unconstrained optimization of the network parameters using standard algorithms while ensuring constraint satisfaction by construction. Furthermore, we show that HardNet retains the universal approximation capabilities of neural networks. We demonstrate the versatility and effectiveness of HardNet across various applications: fitting functions under constraints, learning optimization solvers, optimizing control policies in safety-critical systems, and learning safe decision logic for aircraft systems. [^1] The code is available at https://github.com/azizanlab/hardnet.
@@ 1 Introduction
Neural networks are widely adopted for their generalization capabilities and their ability to model highly non-linear functions in high-dimensional spaces. With their increasing proliferation, it has become more important to be able to impose constraints on neural networks in many applications. By incorporating domain knowledge about input-output relationships into neural networks through constraints, we can enhance their generalization abilities, particularly when the available data is limited (Pathak et al., 2015; Oktay et al., 2017; Raissi et al., 2019). These constraints introduce inductive biases that can guide the model’s learning process toward plausible solutions that adhere to known properties of the problem domain, potentially reducing overfitting to limited data. As a result, neural networks can more effectively capture underlying patterns and make accurate predictions on unseen data, despite the scarcity of training examples.
Moreover, adherence to specific requirements is critical in many practical applications. For instance, in robotics, this could translate to imposing collision avoidance or pose manifold constraints (Ding & Fan, 2014; Wang & Yan, 2023; Ryu et al., 2022; Huang et al., 2022). In geometric learning, this could mean imposing a manifold constraint (Lin & Zha, 2008; Simeonov et al., 2022). In financial risk-management scenarios, violating constraints on the solvency of the portfolio can lead to large fines (McNeil et al., 2015). By enforcing the neural network outputs to satisfy these non-negotiable rules (i.e., hard constraints), we can make the models more reliable, interpretable, and aligned with the underlying problem structure.
However, introducing hard constraints can potentially limit a neural network’s expressive power. To illustrate this point, consider a constraint that requires the neural network’s output to be less than or equal to 1. One could simply restrict the model to always output a constant value less than 1, which ensures the satisfaction of the hard constraint but obviously limits the model capacity drastically. This raises the question:
@@
  • Presenting a practical framework called HardNet (short for hard-constrained neural net) for constructing neural networks that satisfy input-dependent affine/convex constraints by construction. HardNet allows for unconstrained optimization of the networks’ parameters with standard algorithms.
@@
  • Demonstrating the utility of our method on a variety of scenarios where it’s critical to satisfy hard constraints – learning optimization solvers, optimizing control policies in safety-critical systems, and learning safe decision logic for aircraft systems.
@@ 2 Related Work
Early directions focused on implementing data augmentation or domain randomization methods to structure the dataset to satisfy the necessary constraints before training the neural network. However, this does not guarantee constraint satisfaction for arbitrary inputs (especially those far from the training distribution), and the output often violates the constraints marginally on in-distribution inputs as well. Other initial directions focused on introducing the constraints as soft penalties (Márquez-Neila et al., 2017; Dener et al., 2020) to the cost function of the neural network along with Lagrange multipliers as hyperparameters. Raissi et al. (2019); Li et al. (2024) leveraged this idea in their work on physics-informed neural networks (PINNs) to enforce that the output satisfies a given differential equation.
@@
Some of the conventional neural network components can already enforce specific types of hard constraints. For instance, sigmoids can be used for enforcing lower and upper bounds, softmax layers are helpful for imposing simplex constraints, and ReLU layers are essentially projections onto the positive orthant. The convolution layer in ConvNets encodes a translational equivariance constraint which led to significant improvements in empirical performance. Learning new equivariances and inductive biases that accelerate learning for specific tasks is an active area of research.
Recent work has focused on developing new architectures to impose various hard constraints as described in Table 1. Frerix et al. (2020) considered homogeneous linear inequality constraints by embedding a parameterization of the feasible set in a neural network layer. LinSATNet (Wang et al., 2023) enforces positive linear constraints by extending the Sinkhorn algorithm that iteratively finds feasible solutions. Beyond the affine constraints, RAYEN (Tordesillas et al., 2023) imposes a set of convex constraints by parameterizing the feasible set with the neural network output representing a translation from an interior point of the convex feasible set. However, these methods are limited to constraints that depend only on the output (and not the input).
@@
Another line of work considers hard constraints that depend on both input and output. Balestriero & LeCun (2023) proposed the POLICE framework for enforcing the output to be an affine function of the input in certain regions of the input space by reformulating the neural networks as continuous piecewise affine mappings. KKT-hPINN (Chen et al., 2024) enforces more general affine equality constraints by projecting the output to the feasible set where the projection is computed using the KKT conditions of the constraints. However, these affine equality constraints are too restrictive. DC3 (Donti et al., 2021b) is a framework for more general nonlinear constraints which reduces the violations of inequality constraints through gradient-based methods over the manifold where equality constraints are satisfied. However, its constraint satisfaction is not guaranteed in general and is largely affected by the number of gradient steps and the step size, which require fine-tunning.
@@
More closely related to our framework, methods to enforce a single affine inequality constraint are proposed in control literature: Kolter & Manek (2019) presented a framework for learning a stable dynamical model that satisfies a Lyapunov stability constraint. Based on this method, Min et al. (2023) presented the CoILS framework to learn a stabilizing control policy for an unknown control system by enforcing a control Lyapunov stability constraint. Our work generalizes the ideas used in these works to impose more general affine/convex constraints while proving universal approximation guarantees that are absent in prior works; On the theoretical front, Kratsios et al. (2021) presented a constrained universal approximation theorem for probabilistic transformers whose outputs are constrained to be in a feasible set. However, their contribution is primarily theoretical, and they do not present a method for learning such a probabilistic transformer.
Formal Verification of Neural Networks
Verifying whether a provided neural network (after training) always satisfies a set of constraints for a certain set of inputs is a well-studied subject. Albarghouthi et al. (2021) provide a comprehensive summary of the constraint-based and abstraction-based approaches to verification. Constraint-based verifiers are often both sound and complete but they have not scaled to practical neural networks, whereas abstraction-based techniques are approximate verifiers which are sound but often incomplete (Brown et al., 2022; Tjeng et al., 2019; Liu et al., 2021; Fazlyab et al., 2020; Qin et al., 2019; Ehlers, 2017). Other approaches have focused on formally verified exploration and policy learning for reinforcement learning (Bastani et al., 2018; Anderson et al., 2020). Contrary to formal verification methods, which take a pre-trained network and verify that its output always satisfies the desired constraints, our method guarantees constraint satisfaction by construction.
@@
Table 1: Comparison of methods enforcing hard constraints on neural networks for the target function $y=f(x)\in\mathbb{R}^{n_{\text{out}}}$. The baselines above the solid midline consider constraints that only depend on the output. For computation, F and B indicate forward and backward passes, respectively.
@@
Frerix et al. (2020) $Ay\leq 0$ Always F,B: Closed-Form Unknown
@@
RAYEN
(Tordesillas et al., 2023)
$y\in\mathcal{C}$ ($\mathcal{C}$: linear,
quadratic, SOC, LMI)
@@
POLICE
(Balestriero & LeCun, 2023)
$y=Ax+b\;\;\forall x\in R$ Always F,B: Closed-Form Unknown
@@
KKT-hPINN
(Chen et al., 2024)
$Ax+By=b$
(# constraints $\leq n_{\text{out}}$)
Always F,B: Closed-Form Unknown
@@
[…]
F:Iterative
B:Closed-Form
Yes
@@ 3.1 Notation
For $p\in[1,\infty)$, $\|v\|_{p}$ denotes the $\ell^{p}$-norm for a vector $v\in\mathbb{R}^{m}$, and $\|A\|_{p}$ denotes the operator norm for a matrix $A\in\mathbb{R}^{k\times m}$ induced by the $\ell^{p}$-norm, i.e., $\|A\|_{p}=\sup_{w\neq 0}\|Aw\|_{p}/\|w\|_{p}$. $v_{(i)}\in\mathbb{R}$, $v_{(:i)}\in\mathbb{R}^{i}$, and $v_{(i:)}\in\mathbb{R}^{m-i}$ denote the $i$-th component, the first $i$ and the last $m-i$ components of $v$, respectively. Similarly, $A_{(:i)}\in\mathbb{R}^{k\times i}$ and $A_{(i:)}\in\mathbb{R}^{k\times(m-i)}$ denote the first $i$ and the last $m-i$ columns of $A$, respectively.
@@ 3.2 Universal Approximation Theorem
The universal approximation property is a foundational concept in understanding the capabilities of neural networks in various applications. Classical results reveal that shallow neural networks with arbitrary width can approximate any continuous function defined on a compact set as formalized in the following theorem (Cybenko, 1989; Hornik et al., 1989; Pinkus, 1999):
@@
To further understand the success of deep learning, the universal approximation property for deep and narrow neural networks has also been studied in the literature (Lu et al., 2017; Hanin & Sellke, 2017; Kidger & Lyons, 2020; Park et al., 2021). Interesting results show that a critical threshold exists on the width of deep networks that attain the universal approximation property. For instance, deep networks with ReLU activation function with a certain minimum width can approximate any $L^{p}$ function as described in the following theorem (Park et al., 2021, Thm. 1):
@@
For any $p\in[1,\infty)$, $w$-width neural networks with ReLU activation function universally approximate $L^{p}(\mathbb{R}^{n_{\textup{in}}},\mathbb{R}^{n_{\textup{out}}})$ if and only if $w\geq\max\{n_{\textup{in}}+1,n_{\textup{out}}\}$.
@@
Despite these powerful approximation guarantees, they fall short in scenarios where neural networks are required to satisfy hard constraints, such as physical laws or safety requirements. These theorems ensure that a neural network can approximate a target function arbitrarily closely but do not guarantee that the approximation will adhere to necessary constraints. Consequently, even if the target function inherently satisfies specific hard constraints, the neural network approximator might violate them–especially in regions where the target function barely meets the constraints. This shortcoming is particularly problematic for applications that demand strict compliance with non-negotiable domain-specific rules. Therefore, ensuring that neural networks can both approximate target functions accurately and rigorously satisfy hard constraints remains a critical challenge for their deployment in practical applications.
4 HardNet: Hard-Constrained Neural Networks
Figure 1: Illustration of input-dependent constraints and projections performed by HardNet (left) and its schematic diagram (right). A target function $f\!:\!\mathbb{R}\!\rightarrow\!\mathbb{R}^{2}$ satisfies input-dependent hard constraints $f(x)\!\in\!\mathcal{C}(x)$ for each $x\!\in\!\mathbb{R}$. The feasible set $\mathcal{C}(x)$ is visualized as the gray area for two sample inputs $x_{1}$ and $x_{2}$. While the (neural network) function $f_{\theta}$ closely approximates $f$, it violates the hard constraints. Our framework HardNet projects the violated output onto the feasible set in two directions; one in parallel to the boundaries of the satisfied affine constraints (HardNet-Aff) and the other towards the feasible point with the minimum $\ell^{2}$-norm distance from the violated output (HardNet-Cvx). These projections are differentiable, allowing unconstrained optimization of the network parameters using standard algorithms.
@@ 4 HardNet: Hard-Constrained Neural Networks
In this section, we present a practical framework HardNet for enforcing hard constraints on neural networks while retaining their universal approximation properties. In a nutshell, for a parameterized (neural network) function $f_{\theta}:\mathcal{X}\subset\mathbb{R}^{n_{\text{in}}}\rightarrow\mathbb{R}^{n_{\text{out}}}$, we ensure the satisfaction of given constraints by appending a differentiable projection layer $\mathcal{P}$ to $f_{\theta}$. This results in the projected function $\mathcal{P}(f_{\theta}):\mathcal{X}\rightarrow\mathbb{R}^{n_{\text{out}}}$ meeting the required constraints while allowing its output to be backpropagated through to train the model via gradient-based algorithms. Importantly, we show that the proposed architecture has universal approximation guarantees, i.e., it universally approximates the class of functions that satisfy the constraints.
We begin with the simple intuitive case of a single affine constraint and then generalize the approach in two directions. First, we propose HardNet-Aff that ensures compliance with multiple input-dependent affine constraints through a differentiable closed-form projection. Then, we present HardNet-Cvx as a framework to satisfy general input-dependent convex constraints exploiting differentiable convex optimization solvers.
4.1 HardNet-Aff: Imposing Input-Dependent Affine Constraints
First, consider the following single input-dependent affine constraint for a function $f:\mathcal{X}\rightarrow\mathbb{R}^{n_{\text{out}}}$:
$a(x)^{\top}f(x)\leq b(x)\;\;\forall x\in\mathcal{X},$ (1)
where $a(x)\in\mathbb{R}^{n_{\text{out}}}$ and $b(x)\in\mathbb{R}$. We assume $a(x)\neq 0$ as the constraint solely depending on the input is irrelevant to the function. We can ensure this constraint on $\mathcal{P}(f_{\theta})$ by constructing the projection for all $x\in\mathcal{X}$ as:
@@ 4.1 HardNet-Aff: Imposing Input-Dependent Affine Constraints
This closed-form solution is differentiable almost everywhere so that the projected function can be trained using conventional gradient-based algorithms such as SGD. This type of closed-form projection has been recently utilized in the control literature as in Kolter & Manek (2019) for learning stable dynamics and in Donti et al. (2021a); Min et al. (2023) for learning stabilizing controllers. Nonetheless, they are limited to enforcing only a single inequality constraint. Moreover, their empirical success in learning the desired functions has not been theoretically understood. To that end, we generalize the method to satisfy more general constraints and provide a rigorous explanation for its expressivity through universal approximation guarantees.

### v2

@@ Abstract
Incorporating prior knowledge or specifications of input-output relationships into machine learning models has gained significant attention, as it enhances generalization from limited data and leads to conforming outputs. However, most existing approaches use soft constraints by penalizing violations through regularization, which offers no guarantee of constraint satisfaction—an essential requirement in safety-critical applications. On the other hand, imposing hard constraints on neural networks may hinder their representational power, adversely affecting performance. To address this, we propose HardNet, a practical framework for constructing neural networks that inherently satisfy hard constraints without sacrificing model capacity. Unlike approaches that modify outputs only at inference time, HardNet enables end-to-end training with hard constraint guarantees, leading to improved performance. To the best of our knowledge, HardNet is the first method with an efficient forward pass to enforce more than one input-dependent inequality constraint. It allows unconstrained optimization of the network parameters using standard algorithms by appending a differentiable closed-form enforcement layer to the network’s output. Furthermore, we show that HardNet retains the universal approximation capabilities of neural networks. We demonstrate the versatility and effectiveness of HardNet across various applications: learning with piecewise constraints, learning optimization solvers, optimizing control policies in safety-critical systems, and learning safe decision logic for aircraft systems. [^1] The code is available at https://github.com/azizanlab/hardnet.
@@ 1 Introduction
Neural networks are widely adopted for their generalization capabilities and their ability to model highly non-linear functions in high-dimensional spaces. With their increasing proliferation, it has become more important to be able to impose constraints on neural networks in many applications. By incorporating domain knowledge about input-output relationships through constraints, we can enhance generalization, particularly when available data is limited [48, 46, 52]. These constraints introduce inductive biases that guide the model’s learning process toward plausible solutions that adhere to known properties of the problem domain, potentially reducing overfitting. Consequently, neural networks can more effectively capture underlying patterns and make accurate predictions on unseen data, despite scarce training samples.
Moreover, adherence to specific requirements is critical in many practical applications. For instance, in robotics, this could translate to imposing collision avoidance or ensuring configurations remain within a valid motion manifold [18, 67, 53, 31]. In geometric learning, this could mean imposing a manifold constraint [39, 55]. In financial risk management, violating constraints on the solvency of the portfolio can lead to large fines [44]. Enforcing neural network outputs to satisfy these non-negotiable rules (i.e., hard constraints) makes models more reliable, interpretable, and aligned with the underlying problem structure.
However, introducing hard constraints can potentially limit a neural network’s expressive power. To illustrate this point, consider a constraint that requires the model’s output to be less than 1. One could simply restrict the model to always output a constant value less than 1, which ensures the constraint satisfaction but obviously limits the model capacity drastically. This raises the question:
@@
  • Presenting a practical framework called HardNet (short for hard-constrained neural network) for constructing neural networks that satisfy input-dependent constraints by construction. HardNet is, to the best of our knowledge, the first method with an efficient forward pass to enforce more than one input-dependent inequality constraint. It allows for unconstrained optimization of the networks’ parameters with standard algorithms.
@@
  • Demonstrating the utility of our method on a variety of scenarios where it is critical to satisfy hard constraints — learning with piecewise constraints, learning optimization solvers, optimizing control policies in safety-critical systems, and learning safe decision logic for aircraft systems.
@@ 2 Related Work
Early approaches used data augmentation or domain randomization to structure the dataset to satisfy the necessary constraints before training the neural network. However, this does not ensure constraint satisfaction for arbitrary inputs, especially those far from the training distribution, and outputs often violate constraints marginally even on in-distribution inputs. Other initial directions focused on introducing the constraints as soft penalties [43, 17] to the cost function of the neural network along with Lagrange multipliers as hyperparameters. Raissi et al. [52], Li et al. [38] leveraged this idea in their work on physics-informed neural networks (PINNs) to enforce that the output satisfies a given differential equation.
@@
Some conventional neural network components can already enforce specific types of hard constraints. For instance, sigmoids can impose lower and upper bounds, softmax layers enforce simplex constraints, and ReLU layers are projections onto the positive orthant. The convolution layer in ConvNets encodes a translational equivariance constraint, which led to significant improvements in empirical performance. Learning new equivariances and inductive biases that accelerate learning for specific tasks is an active research area.
Recent work has explored new architectures to (asymptotically) impose various hard constraints by either finding certain parameterizations of feasible sets or incorporating differentiable projections into neural networks, as summarized in Table 1. Frerix et al. [25] addressed homogeneous linear inequality constraints by embedding a parameterization of the feasible set in a neural network layer. Huang et al. [30] and LinSATNet [66] introduced differentiable projection methods that iteratively refine outputs to satisfy linear constraints. However, these iterative approaches do not guarantee constraint satisfaction within a fixed number of iterations, limiting their reliability in practice. C-DGM [58] enforces linear inequality constraints in generative models for tabular data by incrementally adjusting each output component in a finite number of iterations. However, its application to input-dependent constraints is limited as it cannot efficiently handle batched data. When constraints are input-dependent, the method requires recomputing the reduced constraint sets for each input, making it computationally prohibitive. In the context of optimal power flow, Chen et al. [15] enforces feasibility for learning optimization proxies through closed-form differentiable repair layers. While effective, this approach is restricted to specific affine constraints.
Beyond affine constraints, RAYEN [63] and Konstantinov & Utkin [35] enforce certain convex constraints by parameterizing the feasible set such that the neural network output represents a translation from an interior point of the convex feasible region. However, these methods are limited to constraints dependent only on the output, and not the input. Extending these methods to input-dependent constraints is challenging because it requires finding different parameterizations for each input, such as determining a new interior point for every feasible set.
@@
Another line of work considers hard constraints that depend on both input and output. Balestriero & LeCun [8] proposed the POLICE framework for enforcing the output to be an affine function of the input in certain regions of the input space by reformulating the neural networks as continuous piecewise affine mappings. KKT-hPINN [14] enforces more general affine equality constraints by projecting the output to the feasible set where the projection is computed using the KKT conditions of the constraints. ACnet [11] enforces nonlinear equality constraints by transforming the nonlinear constraints to affine constraints for redefined inputs and outputs. However, these methods are limited to enforcing only equality constraints. DC3 [20] is a framework for more general nonlinear constraints that reduces the violations of inequality constraints through gradient-based methods over the manifold where equality constraints are satisfied. However, its constraint satisfaction is not guaranteed in general and is largely affected by the number of gradient steps and the step size, which require fine-tuning.
@@
More closely related to our framework, methods to enforce a single affine inequality constraint have been proposed in the control literature: Kolter & Manek [34] presented a framework for learning a stable dynamical model that satisfies a Lyapunov stability constraint. Based on this method, Min et al. [45] presented the CoILS framework to learn a stabilizing control policy for an unknown control system by enforcing a control Lyapunov stability constraint. Our work generalizes the ideas used in these works to impose more general affine/convex constraints while proving universal approximation guarantees that are absent in prior works; On the theoretical front, Kratsios et al. [36] presented a constrained universal approximation theorem for probabilistic transformers whose outputs are constrained to be in a feasible set. However, their contribution is primarily theoretical, and they do not present a method for learning such a probabilistic transformer.
@@
Table 1: Comparison of methods enforcing hard constraints on neural networks for the target function $y=f(x)\in\mathbb{R}^{n_{\text{out}}}$. The baselines above the solid midline consider constraints that only depend on the output. For computation, F and B indicate forward and backward passes, respectively. HardNet-Aff is the only method to enforce input-dependent constraints provably and efficiently with universal approximation guarantees.
@@
Frerix et al. [25] $Ay\leq 0$ Always F,B: Closed-Form Unknown
@@
[…]
$y\in\mathcal{C}$ ($\mathcal{C}$: linear, quadratic, SOC, LMI)
@@
Formal Verification of Neural Networks
Verifying whether a provided neural network (after training) always satisfies a set of constraints for a certain set of inputs is a well-studied subject. Albarghouthi et al. [3] provide a comprehensive summary of the constraint-based and abstraction-based approaches to verification. Constraint-based verifiers are often both sound and complete but they have not scaled to practical neural networks, whereas abstraction-based techniques are approximate verifiers which are sound but often incomplete [13, 22, 12, 62, 40, 23, 51, 21]. Other approaches have focused on formally verified exploration and policy learning for reinforcement learning [9, 6, 65]. Contrary to most formal verification methods, which take a pre-trained network and verify that its output always satisfies the desired constraints, our method guarantees constraint satisfaction by construction throughout the training.
To address constraint violations in trained models after verification, some methods in safe reinforcement learning introduce shielding mechanisms by overriding unsafe decisions using a backup policy based on reachability analysis [54, 10]. While effective, this setup introduces a hard separation between learning and enforcement, often sacrificing performance and limiting joint optimization. Shielding mechanisms are typically not differentiable and cannot be integrated into training. Similarly, editing methods modify trained networks post hoc to enforce output constraints by solving relaxed optimization problems over the model parameters [56, 60]. Though recent works demonstrate their application during training, these techniques generally target input-independent constraints and are architecture-dependent.
@@ 3.1 Notation
For $p\in[1,\infty)$, $\|v\|_{p}$ denotes the $\ell^{p}$-norm for a vector $v\in\mathbb{R}^{m}$, and $\|A\|_{p}$ denotes the operator norm for a matrix $A\in\mathbb{R}^{k\times m}$ induced by the $\ell^{p}$-norm, i.e., $\|A\|_{p}=\sup_{w\neq 0}\|Aw\|_{p}/\|w\|_{p}$. $v_{(i)}\in\mathbb{R}$, $v_{(:i)}\in\mathbb{R}^{i}$, and $v_{(i:)}\in\mathbb{R}^{m-i}$ denote the $i$-th component, the first $i$ and the last $m-i$ components of $v$, respectively. Similarly, $A_{(:i)}\in\mathbb{R}^{k\times i}$ and $A_{(i:)}\in\mathbb{R}^{k\times(m-i)}$ denote the first $i$ and the last $m-i$ columns of $A$, respectively. $[A;B]$ denotes the row-wise concatenation of the matrices $A$ and $B$.
@@ 3.2 Universal Approximation Theorem
The universal approximation property is a foundational concept in understanding the capabilities of neural networks in various applications. Classical results reveal that shallow neural networks with arbitrary width can approximate any continuous function defined on a compact set as formalized in the following theorem [16, 29, 37, 49]:
@@
To further understand the success of deep learning, the universal approximation property for deep and narrow neural networks has also been studied in the literature [41, 28, 33, 47]. Interesting results show that a critical threshold exists on the width of deep networks that attain the universal approximation property. For instance, deep networks with ReLU activation function with a certain minimum width can approximate any $L^{p}$ function as described in the following theorem [47, Thm. 1]:
@@
For any $p\in[1,\infty)$, $w$-width neural networks with $\mathsf{ReLU}{}$ activation function universally approximate $L^{p}(\mathbb{R}^{n_{\textup{in}}},\mathbb{R}^{n_{\textup{out}}})$ if and only if $w\geq\max\{n_{\textup{in}}+1,n_{\textup{out}}\}$.
@@
Despite these powerful approximation guarantees, they fall short when neural networks are required to satisfy hard constraints, such as physical laws or safety requirements. These theorems ensure that a neural network can approximate a target function arbitrarily closely but do not guarantee adherence to necessary constraints. Consequently, even if the target function satisfies specific hard constraints, the neural network approximator might violate them–especially in regions where the target function barely meets the constraints. This shortcoming is particularly problematic for applications that demand strict compliance with non-negotiable domain-specific rules. Therefore, ensuring that neural networks can both approximate target functions accurately and rigorously satisfy hard constraints remains a critical challenge for their deployment in practical applications.
Figure 1: Schematic of HardNet. Its differentiable enforcement layer allows unconstrained end-to-end optimization of the network parameters using standard algorithms while guaranteeing satisfaction with input-dependent constraints by construction. The layer can be applied to any neural networks.
4 HardNet: Hard-Constrained Neural Network
@@
In this section, we present a practical framework HardNet shown in Figure 1 for enforcing input-dependent hard constraints on neural networks while retaining their universal approximation properties. In a nutshell, for a parameterized (neural network) function $f_{\theta}:\mathcal{X}\subset\mathbb{R}^{n_{\text{in}}}\rightarrow\mathbb{R}^{n_{\text{out}}}$, we ensure the satisfaction of given constraints by appending a differentiable enforcement layer with a projection $\mathcal{P}$ to $f_{\theta}$. This results in the projected function $\mathcal{P}(f_{\theta}):\mathcal{X}\rightarrow\mathbb{R}^{n_{\text{out}}}$ meeting the required constraints while allowing its output to be backpropagated through to train the model via gradient-based algorithms. Importantly, we show that the proposed architecture has universal approximation guarantees, i.e., it universally approximates the class of functions that satisfy the constraints.
A key challenge in this approach is devising a differentiable projection that has efficient forward and backward passes. For instance, consider affine constraints $A(x)f(x)\leq b(x)\;\forall x\in\mathcal{X}$ for a function $f:\mathcal{X}\rightarrow\mathbb{R}^{n_{\text{out}}}$ with $A(x)\in\mathbb{R}^{n_{\text{ineq}}\times n_{\text{out}}}$ and $b(x)\in\mathbb{R}^{n_{\text{ineq}}}$, one would have
$\mathcal{P}(f_{\theta})(x)=\operatorname*{arg\,min}_{z\in\mathbb{R}^{n_{\text{out}}}}\|z\!-\!f_{\theta}(x)\|_{2}\text{ s.t. }A(x)z\leq b(x).$ (1)
Although the constraints are affine, this projection does not admit a closed-form solution in general, making it expensive to compute–especially as it needs to be computed over every sample $x$.
In case of a single constraint $a(x)^{\top}f(x)\leq b(x)\;\;\forall x\in\mathcal{X}$ with $a(x)\in\mathbb{R}^{n_{\text{out}}}$ and $b(x)\in\mathbb{R}$, recent work in the control literature–for instance, Kolter & Manek [34] for learning stable dynamics and Donti et al. [19], Min et al. [45] for learning stabilizing controllers–has adopted the following closed-form solution:
@@
This solution is differentiable almost everywhere so that the projected function can be trained using standard optimizers.
Example 4.1.
The constraint $(f(x))_{(0)}\!\geq\!x(f(x))_{(1)}$ on $f\!:\!\mathbb{R}\!\rightarrow\!\mathbb{R}^{2}$ can be encoded as $a(x)\!=\![-1;x],b(x)\!=\!0$ so that a sample $f_{\theta}(1)\!=\![3;5]$ is projected to $\mathcal{P}(f_{\theta})(1)\!=\![4;4]$, satisfying the constraint.
Nonetheless, the formulation is limited to enforcing only a single inequality constraint. Moreover, its empirical success in learning the desired functions has not been theoretically understood. To that end, we propose HardNet-Aff, the first method to the best of our knowledge, with an efficient forward pass to enforce more than one input-dependent inequality constraint. In addition, we provide a rigorous explanation for its expressivity through universal approximation guarantees. Then, we discuss HardNet-Cvx as a conceptual framework to satisfy general input-dependent convex constraints exploiting differentiable convex optimization solvers.
4.1 HardNet-Aff: Imposing Input-Dependent Affine Constraints